# MRD: Using Physically Based Differentiable Rendering to Probe Vision Models for 3D Scene Understanding

### v1

@@ 1. Introduction
2. Related work
Synthesis-based explanation methods have a long history in vision science and computer vision. The central idea is to understand the representations of image-computable vision models by synthesizing new images via an optimization process that minimizes some objective function. (Other visualization-based explanation methods primarily highlight pixels in the target image that are somehow correlated with a model’s decision (Simonyan et al., 2014; Selvaraju et al., 2017; Ribeiro et al., 2016), but will not be discussed further here). In many previous applications, this involved pixel-based gradient ascent: starting from white-noise images, and iteratively adjusting the pixels. This approach has been used to generate equivalence classes (metamers) for image-computable vision models, leading to significant impact in both computer vision and human vision science (Balas et al., 2009; Portilla and Simoncelli, [n. d.]; Freeman and Simoncelli, 2011; Wallis et al., 2016; Boehm et al., 2025; Broderick et al., 2023; Walton et al., 2021; Balas, 2006; Rosenholtz et al., 2012; Feather et al., 2023, 2019). It is also possible to synthesize images that efficiently discriminate between competing models (Golan et al., 2020; Wang and Simoncelli, 2008; Berardino et al., 2017; Feather et al., 2025; Boehm et al., 2025). Feature visualization and related techniques (Erhan et al., 2009; Mahendran and Vedaldi, 2015; Olah et al., 2017; Nguyen et al., 2019) are also forms of synthesis-based explanation. Here, images are typically generated to maximize the activation of a network layer or node. These methods are argued to provide qualitative insights into network activations and therefore explanations for their decisions, though the usefulness of these methods in providing humans with unique explanatory information has recently been called into question (Zimmermann et al., 2021; Borowski et al., 2021; Geirhos et al., 2024).
@@ 3. General Methods
In the following, we will briefly recap the methods from the different fields on which our work is based on.
@@ 3.4. Implementation Details
We implement all of our experiments using Python 3.12, PyTorch (Paszke et al., 2019), and Mitsuba 3 (Jakob et al., 2022), which is a suitable PBDR for our experiments, and mitigates many of the discontinuity issues mentioned above ((Nimier-David et al., 2020; Vicini et al., 2021; Nicolet et al., 2021; Zhang et al., 2023, 2024)). For all of our experiments, we also define an early stopping criterion: the experiment stops if the loss does not improve for 50 epochs. We run all of our shape reconstruction experiments on a workstation with a single RTX 4090 with 24 GB of VRAM using Ubuntu 22.04.5 LTS. For our material reconstruction experiments, more VRAM is required due to the path tracing integrator; for these experiments, we resort to an RTX 6000 with 48 GB of VRAM running on an identical system environment. We use pretrained models and implementations of these if available (Geirhos et al., 2019; Radford et al., 2021). Our code will be made available after review.
@@ 3.5. Baseline experiments
Translucent 4 $3\times 10^{-2}$ 500 roughness, $\eta$, albedo texture, spec_trans
@@ 4.2. Results
Brushed Metal ResNet-SIN 110 0.998 0.998 - - 0.000
@@
Translucent CLIP 124 0.706 0.653 -0.067 No 0.054
Translucent DINO 190 0.771 0.710 0.333 No 0.061
Translucent LPIPS 88 0.831 0.810 - - 0.021
Translucent ResNet 191 0.973 0.956 - - 0.016
Translucent ResNet-SIN 94 0.998 0.998 - - 0.000
Translucent VGG 60 0.635 0.600 - - 0.034
@@
Across our different scenes, we were able to find metamers for most of the networks tested (Table 3). Only ResNet and, in one instance, CLIP, had examples that did not qualify as metamers under our definition. Subjective visual results appeared reasonable (Figure 3 shows example reconstructions of the brushed metal materials for VGG and ResNet).
@@
Table 3 summarizes the per-BSDF / per-network outcomes. The metameric results attain a top similarity that meets or exceeds the baseline (similarity difference $\gtrsim 0$), indicating that the reconstructed BSDF is representationally indistinguishable from the ground truth under the evaluated representation. In the case of non-metamers, the top similarity is noticeably lower than the baseline (negative similarity difference), indicating that the reconstruction did not satisfy the material metamer criterion for that network. Overall, we find 21 metamers and four non-metamers.
@@
Different material classes show systematically different recoverability. Metallic and specular BSDFs (brushed-metal, aurora-like materials) tend to produce strong global image cues (high-contrast highlights, anisotropic streaks) that generate coherent gradients and are thus easier to recover to baseline similarity. Diffuse and strongly translucent materials can be more challenging: they sometimes produce broader similarity gaps or lower RSA values, likely because their appearance depends on subtle indirect-lighting interactions or transmission effects that are harder to disambiguate from other scene factors. Where translucency interacts with geometry and indirect illumination, the optimization needs a larger number of samples and iterations to reach the same representational fidelity.
@@ 5.2. Results
Figure 5 demonstrates the convergence behavior and resulting metameric geometries for two of our experiments (Dragon Hallstatt using LPIPS as the metric / latent space, and Dragon Garden with ResNet as the latent). The reconstruction similarity consistently improves, converging towards the baseline performance (orange curve). We specifically analyze the iteration where the reconstruction fidelity peaks or intersects with the baseline threshold. In the Dragon Hallstatt LPIPS experiment, the reconstruction similarity exceeds the baseline similarity (making this a metamer according to our definition). The visual output at this intersection point is subjectively indistinguishable from the baseline. This confirms that MRD can successfully recover scene parameters that appear equivalent to the target under the LPIPS representation.
However, visual results for the Dragon Garden ResNet experiment are strikingly different. First, the reconstruction similarity does not exceed the baseline similarity, making this non-metameric by our strict definition. However, the two similarities are quite close, and in addition the RSA correlation of 0.67 is significant (see Table 4). Visually, however, the ResNet results look quite unlike a dragon to a human viewer. That the similarities and RSA scores for ResNet are high despite these qualitative differences is consistent with the notion that ResNet has a very wide equivalence class for 3D shapes: spiky blobs look like dragons to ResNet. We observed similar results for ResNet applied to other geometries.
@@
Table 4. Shape reconstruction results across all shapes, environment maps and NNs. The table shows the epoch of the top recorded similarity (averaged across views) and compares it to the top similarity measured during base reconstruction. If the NN produces a latent space we further measure the RSA and its significance values. Lastly, the similarity difference is provided to see the marginal errors.
Experiment (Shape + Envmap + Model) Epoch Top Sim Top Base RSA Significant ($p<0.05$) Similarity Difference
METAMERS Dog Garden LPIPS 159 0.991 0.983 - - 0.008
Dog Garden VGG 164 0.983 0.963 - - 0.020
Dog Hallstatt LPIPS 158 0.991 0.991 - - 0.000
Dog Hallstatt VGG 258 0.982 0.979 - - 0.003
Dog Skybox VGG 499 0.980 0.835 - - 0.144
Dragon Skybox LPIPS 248 0.798 0.782 - - 0.016
Dragon Skybox ResNet 238 0.953 0.945 0.536 Yes 0.008
Dragon Skybox VGG 486 0.578 0.519 - - 0.058
Lion Statue Hallstatt LPIPS 348 0.957 0.956 - - 0.001
Suzanne Garden LPIPS 115 0.986 0.972 - - 0.015
Suzanne Garden VGG 118 0.970 0.937 - - 0.034
Suzanne Hallstatt LPIPS 159 0.985 0.979 - - 0.006
Suzanne Hallstatt VGG 101 0.968 0.953 - - 0.015
Suzanne Skybox LPIPS 236 0.953 0.933 - - 0.020
Suzanne Skybox VGG 158 0.872 0.849 - - 0.023
@@
Figure 6. Empirical cumulative distributions of LPIPS (Dragon Hallstatt) for three conditions: the null distribution (orange) obtained from random, unrelated objects; the seeds distribution (teal) showing variability across independent reconstruction runs with different initializations; and the baseline distribution (purple), representing the similarity achieved during the ground-truth optimization run used for hyperparameter selection. Tick marks along the x-axis indicate individual samples. Higher similarity values denote closer alignment between latent representations.
@@
As a control experiment to verify our criterion for metamerism, we compute the empirical cumulative distribution function (ECDF) for three shape conditions: (1) a null distribution, being the hyperspherical similarity between the ground truth target mesh renderings and 1000 random object meshes, (2) multiple reconstructions with different seeds, and (3) similarity measured during ground truth run. An example ECDF for LPIPS on the Dragon-Hallstatt scene (Figure 6) reveals a clear separation between the conditions. The null distribution occupies the lowest range of similarities, confirming that unrelated objects produce widely scattered and generally low latent correlations. In contrast, the seeds distribution clusters tightly near high similarity values, indicating that independent reconstruction runs converge to nearly identical similarity values despite differing initial seeds. Finally, the baseline distribution lies at the extreme upper end of the similarity spectrum, reflecting the latent alignment achieved during the ground-truth run. This hierarchical ordering, $null<seeds\approx baseline$, supports our metamer criterion: reconstructions qualify as model metamers when their latent similarity matches the characteristic similarity levels measured in the baseline condition.
@@ 6. Discussion
Across all material experiments, MRD achieved metamerism in a substantial subset of scenes, with many networks reaching or closely approaching their ground truth baseline similarities (Table 3). The LPIPS perceptual metric was particularly effective for materials, consistently producing metamers across nearly all tested BSDF classes (Aurora, Brushed Metal, Diffuse, Translucent) with top similarity scores frequently exceeding $0.95$ and minimal deviations from the baseline run. Similarly, VGG and CLIP demonstrated strong performance for several materials, though CLIP exhibited higher variability: certain material–lighting conditions achieved exceptionally high RSA values (e.g., Aurora CLIP, RSA =1.0=1.0) despite moderately lower absolute similarity scores. Although a portion of the BSDF experiments did not strictly meet the metamer criterion, many of these failures were marginal. Networks such as ResNet-SIN often produced reconstructions with extremely small similarity gaps (e.g., $\Delta_{\text{sim}}<0.02$ for brushed metal and translucent materials), indicating that the optimized BSDFs were nearly representationally identical to the ground truth. These marginal failures suggest that small amounts of optimization noise or local irregularities in the BSDF parameter landscape, rather than representational mismatch, prevented some reconstructions from formally qualifying as metamers. Overall, material reconstruction appears robust across networks, with LPIPS and VGG providing the highest similarities.
@@
Shape reconstruction was more challenging overall, with far fewer scenes meeting our metamer threshold (Table 4). While LPIPS remained a strong metric, achieving metamers across several of the evaluated object classes (Dog, Dragon, Lion Statue, Suzanne), shape recovery generally required more iterations and exhibited greater network-dependent variability. VGG also succeeded in several cases (e.g., Suzanne Garden VGG), reflecting its sensitivity to mid-level image structure that is strongly influenced by shape. However, CLIP and ResNet variants showed more variable results for geometry: in many conditions, similarity fell just short of baseline despite the reconstructions visually matching the target geometry. These marginal failures are particularly prominent for ResNet and ResNet-SIN, where several experiments achieved exceptionally high absolute similarity (often >0.97>0.97) but missed the metameric threshold by only a small margin (e.g., $\Delta_{\text{sim}}\approx 0.016$$0.025$ for Dragon scenes). Such small differences indicate that although the reconstructed shapes did not strictly surpass the baseline similarity, their latent representations were almost identical to the ground truth. These results suggest that shape reconstruction is more susceptible to optimization noise and local minima in the high-dimensional geometry space, but that the method nonetheless produces high-quality latents that are very close to metameric for several networks, even though the visual reconstructions often yield “anamorphous blobs" (as seen in Figure 5).
@@
We observe consistently higher reconstruction similarity and more stable convergence for material (BSDF) reconstruction compared to shape reconstruction. This difference can be explained by several factors. (1) Material variation occupies a substantially lower-dimensional and more structured subspace than shape. Changes in BRDF parameters such as roughness, metallicity, or anisotropy produce smooth, predictable modifications in shading and reflections, which map more linearly into the model’s latent space. In contrast, shape variation is inherently high-dimensional: small geometric perturbations can induce large, non-linear changes in silhouette, occlusion, and local shading, making the latent response more complex and less predictable. (2) Material edits affect the image globally, providing spatially dense and coherent gradients during optimization, whereas shape edits often result in sparse or discontinuous gradients that slow convergence and increase the likelihood of suboptimal solutions. (4) Modern vision networks have been shown to encode texture and shading statistics more explicitly than geometric information (Geirhos et al., 2019); as a result, latent embeddings may be naturally more sensitive to material differences than to shape differences. (5) The optimization landscape for materials is smoother, with fewer local minima, and the ground truth material embeddings exhibit lower intra-class variance. Together, these factors make it significantly easier for the optimizer to match the baseline similarity for materials, leading to improved reconstruction quality and more reliable identification of material metamers.
@@
Recently, Elumalai et al. (Elumalai et al., 2025) optimized a mesh to obtain maximally exciting inputs for a Robust ResNet’s activation using a differentiable rasterizer. While this approach is broadly similar to our work, in that it uses gradient-based differentiable rendering to optimize against a network activation objective, rasterization inherently lacks physical light transport and (for example) cannot capture global illumination or material interactions beyond direct shading. In contrast, MRD reconstructs latent representations using a differentiable path tracer, thereby accounting for physically based light transport and simulating light propagation in a 3D scene.

### v3

@@ 1. Introduction
1.1. Related work
Synthesis-based explanation methods have a long history in vision science and computer vision. The central idea is to understand the representations of image-computable vision models by synthesizing new images via an optimization process that minimizes some objective function. (Other visualization-based explanation methods primarily highlight pixels in the target image that are somehow correlated with a model’s decision (Simonyan et al., 2014; Selvaraju et al., 2017; Ribeiro et al., 2016), but will not be discussed further here). In many previous applications, this involved pixel-based gradient ascent: starting from white-noise images, and iteratively adjusting the pixels. This approach has been used to generate equivalence classes (metamers) for image-computable vision models, leading to significant impact in both computer vision and human vision science (Balas et al., 2009; Portilla and Simoncelli, [n. d.]; Freeman and Simoncelli, 2011; Wallis et al., 2016; Boehm et al., 2025; Broderick et al., 2023; Walton et al., 2021; Balas, 2006; Rosenholtz et al., 2012; Feather et al., 2023, 2019). It is also possible to synthesize images that efficiently discriminate between competing models (Golan et al., 2020; Wang and Simoncelli, 2008; Berardino et al., 2017; Feather et al., 2025; Boehm et al., 2025). Feature visualization and related techniques (Erhan et al., 2009; Mahendran and Vedaldi, 2015; Olah et al., 2017; Nguyen et al., 2019) are also forms of synthesis-based explanation. Here, images are typically generated to maximize the activation of a network layer or node. These methods are argued to provide qualitative insights into network activations and therefore explanations for their decisions, though the usefulness of these methods in providing humans with unique explanatory information has recently been called into question (Zimmermann et al., 2021; Borowski et al., 2021; Geirhos et al., 2024). Another approach has been proposed by Veeravasarapu et al. (Veeravasarapu et al., 2017) by training a GAN using a Bayesian generative model receiving a set of scene parameters and a CNN as a discrimnator to optimize the generative model via Bayesian updates to produce 3D scenes in a probabilistic manner.
@@ 2. General Methods
In the following, we will briefly recap the methods from the different fields on which our work is based.
@@ 2.4. Implementation Details
We implement all of our experiments using Python 3.12, PyTorch (Paszke et al., 2019), and Mitsuba 3 (Jakob et al., 2022), which is a suitable PBDR for our experiments, and mitigates many of the discontinuity issues mentioned above ((Nimier-David et al., 2020; Vicini et al., 2021; Nicolet et al., 2021; Zhang et al., 2023, 2024)). For all of our experiments, we also define an early stopping criterion: the experiment stops if the loss does not improve for 50 epochs. We run all of our shape reconstruction experiments on a workstation with a single RTX 4090 with 24 GB of VRAM using Ubuntu 22.04.5 LTS. For our material reconstruction experiments, more VRAM is required due to the path tracing integrator; for these experiments, we resort to an RTX 6000 with 48 GB of VRAM running on an identical system environment. We use pretrained models and implementations of these if available (Geirhos et al., 2019; Radford et al., 2021). Our code will be available here: https://github.com/ag-perception-wallis-lab/MRD.
@@ 2.5. Baseline experiments
Translucent 4 $2\times 10^{-1}$ 500 roughness, $\eta$, albedo texture, spec_trans
@@ 3.2. Results
Brushed Metal ResNet-SIN 110 0.998 0.998 0.333 No 0.000
@@
[…]
Across our different scenes, we were able to find metamers for most of the networks tested (Table 3). The translucent reconstructions, ResNet and CLIP, had examples that did not qualify as metamers under our definition. Subjective visual results appeared reasonable (Figure 3 shows example reconstructions of the brushed metal materials for VGG and ResNet).
@@
Table 3 summarizes the per-BSDF / per-network outcomes. The metameric results attain a top similarity that meets or exceeds the baseline (similarity difference $\gtrsim 0$), indicating that the reconstructed BSDF is representationally indistinguishable from the ground truth under the evaluated representation. In the case of non-metamers, the top similarity is noticeably lower than the baseline (negative similarity difference), indicating that the reconstruction did not satisfy the material metamer criterion for that network. Overall, we find 14 metamers and 10 non-metamers.
@@
Different material classes show systematically different recoverability. Metallic and specular BSDFs (brushed-metal, aurora-like materials) tend to produce strong global image cues (high-contrast highlights, anisotropic streaks) that generate coherent gradients and are thus easier to recover to baseline similarity. Diffuse and translucent materials can be more challenging: they sometimes produce broader similarity gaps or lower RSA values, likely because their appearance depends on subtle indirect-lighting interactions or transmission effects that are harder to disambiguate from other scene factors. Where translucency interacts with geometry and indirect illumination, the optimization needs a larger number of samples and iterations to reach the same representational fidelity.
@@ 4.2. Results
Figure 5 demonstrates the convergence behavior and resulting metameric geometries for two of our experiments (Dragon Hallstatt using LPIPS as the metric / latent space, and Dragon Garden with ResNet as the latent). The reconstruction similarity consistently improves, converging towards the baseline performance (orange curve). We specifically analyze the iteration where the reconstruction fidelity peaks or intersects with the baseline threshold. In the Dragon Hallstatt LPIPS experiment, the reconstruction similarity exceeds the baseline similarity (making this a metamer according to our definition). The visual output at this intersection point is subjectively indistinguishable from the baseline. However, visual results for the Dragon Garden ResNet experiment are strikingly different. First, the reconstruction similarity does not exceed the baseline similarity, making this non-metameric by our strict definition. However, the two similarities are quite close, and in addition the RSA correlation of 0.67 is significant (see Table 4). Visually, however, the ResNet results look quite unlike a dragon to a human viewer. That the similarities and RSA scores for ResNet are high despite these qualitative differences is consistent with the notion that ResNet has a very wide equivalence class for 3D shapes: spiky blobs look like dragons to ResNet. We observed similar results for ResNet applied to other geometries.
@@
Table 4. Shape reconstruction results across all shapes, environment maps and NNs. The table shows the epoch of the top recorded similarity (averaged across views) and compares it to the top similarity measured during base reconstruction. If the NN produces a latent space we further measure the RSA and its significance values. Furthermore, the similarity difference is provided to see the marginal errors. Lastly, we compute the difference between seeds and baseline/null ECDFs (see Figure 6, approximate a Gaussian kernel density and normalizing the $x$-range to [1,1][-1,1]. The vertical bars show [1,0,1][-1,0,1] from left to right. This table is best viewed digitally.
Experiment (Shape + Envmap + Model) Epoch Top Sim Top Base RSA Significant ($p<0.05$) Similarity Difference Seed Null Diff Seed Base Diff
METAMERS Dog Garden LPIPS 159 0.991 0.983 - - 0.008 [Uncaptioned image] [Uncaptioned image]
Dog Garden VGG 164 0.983 0.963 - - 0.020 [Uncaptioned image] [Uncaptioned image]
Dog Hallstatt LPIPS 158 0.991 0.991 - - 0.000 [Uncaptioned image] [Uncaptioned image]
Dog Hallstatt VGG 258 0.982 0.979 - - 0.003 [Uncaptioned image] [Uncaptioned image]
Dog Skybox VGG 499 0.980 0.835 - - 0.144 [Uncaptioned image] [Uncaptioned image]
Dragon Skybox LPIPS 248 0.798 0.782 - - 0.016 [Uncaptioned image] [Uncaptioned image]
Dragon Skybox ResNet 238 0.953 0.945 0.536 Yes 0.008 [Uncaptioned image] [Uncaptioned image]
Dragon Skybox VGG 486 0.578 0.519 - - 0.058 [Uncaptioned image] [Uncaptioned image]
Lion Statue Hallstatt LPIPS 348 0.957 0.956 - - 0.001 [Uncaptioned image] [Uncaptioned image]
Suzanne Garden LPIPS 115 0.986 0.972 - - 0.015 [Uncaptioned image] [Uncaptioned image]
Suzanne Garden VGG 118 0.970 0.937 - - 0.034 [Uncaptioned image] [Uncaptioned image]
Suzanne Hallstatt LPIPS 159 0.985 0.979 - - 0.006 [Uncaptioned image] [Uncaptioned image]
Suzanne Hallstatt VGG 101 0.968 0.953 - - 0.015 [Uncaptioned image] [Uncaptioned image]
Suzanne Skybox LPIPS 236 0.953 0.933 - - 0.020 [Uncaptioned image] [Uncaptioned image]
Suzanne Skybox VGG 158 0.872 0.849 - - 0.023 [Uncaptioned image] [Uncaptioned image]
@@
Figure 6. The mean empirical cumulative distributions of across all experiments (three environment maps, four geometries and six models). We show three distributions. (a) The null distribution (orange) obtained from random, unrelated objects; (b) the seeds distribution (teal) showing variability across independent reconstruction runs with different initializations; (c) and the baseline distribution (purple), representing the similarity achieved during the ground-truth optimization run used for hyperparameter selection. Higher similarity values denote closer alignment between latent representations. The shaded region is one standard deviation.
@@
As a control experiment to verify our criterion for metamerism in the shape experiments, we compute the empirical cumulative distribution function (ECDF) for three shape conditions: (1) a null distribution, being the hyperspherical similarity between the ground truth target mesh renderings and 1000 random object meshes, (2) multiple reconstructions with different seeds, and (3) similarity measured during ground truth run. We show the average ECDF across all shape reconstructions in Figure 6, showing a clear separation of the null and the other distributions. Furthermore, we plot the differences between the null and seed distribution, as well as the difference between null and baseline distribution in Table 4. The null distribution occupies the lowest range of similarities, confirming that unrelated objects produce widely scattered and generally low latent correlations. In contrast, the seeds distribution clusters tightly near high similarity values, indicating that independent reconstruction runs converge to nearly identical similarity values despite differing initial seeds. Finally, the baseline distribution lies at the extreme upper end of the similarity spectrum, reflecting the latent alignment achieved during the ground-truth run. This hierarchical ordering, $null<seeds\approx baseline$, supports our metamer criterion: reconstructions qualify as model metamers when their latent similarity matches the characteristic similarity levels measured in the baseline condition. If network representations simply found any mesh to bi similar to the target, then we would expect the distribution of similarities for random meshes to overlap with that for the reconstructions (seeds). Instead, we find that the models show variable behavior where VGG similarity values are more or less between [0.777,0.020][-0.777,0.020] and ResNet-SIN with very narrow similarities between $[0.997,0.998]$.
@@ 5. Discussion
Across all material experiments, MRD achieved metamerism in a substantial subset of scenes, with many networks reaching or closely approaching their baseline similarities (Table 3). The LPIPS perceptual metric was particularly effective for materials, consistently producing metamers across nearly all tested BSDF classes (Aurora, Brushed Metal, Diffuse, Translucent) with top similarity scores frequently exceeding $0.95$ and minimal deviations from the baseline run. Similarly, VGG and CLIP demonstrated strong performance for several materials, though CLIP exhibited higher variability: certain material–lighting conditions achieved exceptionally high RSA values (e.g., Aurora CLIP, RSA =1.0=1.0) despite moderately lower absolute similarity scores. Although a portion of the BSDF experiments did not strictly meet the metamer criterion, many of these failures were marginal. Networks such as ResNet-SIN often produced reconstructions with extremely small similarity gaps (e.g., $\Delta_{\text{sim}}<0.02$ for brushed metal and translucent materials), indicating that the optimized BSDFs were nearly representationally identical to the baseline. These marginal failures suggest that small amounts of optimization noise or local irregularities in the BSDF parameter landscape, rather than representational mismatch, prevented some reconstructions from formally qualifying as metamers. Overall, material reconstruction appears robust across networks, with LPIPS and VGG providing the highest similarities.
@@
Shape reconstruction was more challenging overall, with far fewer scenes meeting our metamer threshold (Table 4). While LPIPS remained a strong metric, achieving metamers across several of the evaluated object classes (Dog, Dragon, Lion Statue, Suzanne), shape recovery generally required more iterations and exhibited greater network-dependent variability. VGG also succeeded in several cases (e.g., Suzanne Garden VGG), reflecting its sensitivity to mid-level image structure that is strongly influenced by shape. However, CLIP and ResNet variants showed more variable results for geometry: in many conditions, similarity fell just short of baseline despite the reconstructions visually matching the target geometry. These marginal failures are particularly prominent for ResNet and ResNet-SIN, where several experiments achieved exceptionally high absolute similarity (often >0.97>0.97) but missed the metameric threshold by only a small margin (e.g., $\Delta_{\text{sim}}\approx 0.016$$0.025$ for Dragon scenes). Such small differences indicate that although the reconstructed shapes did not strictly surpass the baseline similarity, their latent representations were almost identical to the baseline. These results suggest that shape reconstruction is more susceptible to optimization noise and local minima in the high-dimensional geometry space, but that the method nonetheless produces high-quality latents that are very close to metameric for several networks, even though the visual reconstructions often yield “anamorphous blobs" (as seen in Figure 5).
@@
We observe consistently higher reconstruction similarity and more stable convergence for material (BSDF) reconstruction compared to shape reconstruction. This difference can be explained by several factors. (1) Material variation occupies a substantially lower-dimensional and more structured subspace than shape. Changes in BRDF parameters such as roughness, metallicity, or anisotropy produce smooth, predictable modifications in shading and reflections, which map more linearly into the model’s latent space. In contrast, shape variation is inherently high-dimensional: small geometric perturbations can induce large, non-linear changes in silhouette, occlusion, and local shading, making the latent response more complex and less predictable. (2) Material edits affect the image globally, providing spatially dense and coherent gradients during optimization, whereas shape edits often result in sparse or discontinuous gradients that slow convergence and increase the likelihood of suboptimal solutions. (4) Modern vision networks have been shown to encode texture and shading statistics more explicitly than geometric information (Geirhos et al., 2019); as a result, latent embeddings may be naturally more sensitive to material differences than to shape differences. (5) The optimization landscape for materials is smoother, with fewer local minima, and the baseline material embeddings exhibit lower intra-class variance. Together, these factors make it significantly easier for the optimizer to match the baseline similarity for materials, leading to improved reconstruction quality and more reliable identification of material metamers.
@@
Recently, Elumalai et al. (Elumalai et al., 2025) optimized a mesh to obtain maximally exciting inputs for a Robust ResNet’s activation using a differentiable rasterizer reconstucting. In their work, they successfully use differentiable rendering to optimize against a network activation objective, yet, rasterization lacks physicality as it is bound to approximations of global illumination and/or screen-space techniques. In contrast, MRD reconstructs latent representations using a differentiable path tracer, thereby accounting for physically based light transport and simulating light propagation in a 3D scene.